Brenda is driving in her new car on her way to class. She is approaching a dip in the road, where the radius of the dip is 28.4 meters. What must the speed of her vehicle be at the bottom of the dip so that she feels three times her weight? (answer is 23.6 m/s)
Answer #1

v = sqrt(2 g h) = sqrt((2)*(9.8)*(28.4)) = 23.6 m/s